Marc Horowitz 21.512

Reactions to “Illness as Metaphor”

In Susan Sontag’s essay, “Illness as Metaphor”, she gives many examples of how cultures as old as Ancient Rome have used illnesses as metaphors for societal issues and problems. My primary reaction is that of a scientist. Namely, that these societies in many cases placed meanings in disease which were often anti-causal, or unrelated. Society creates an image that a weak “life force” brings about tuberculosis, or that living in a rigid manner causes cancer, while in the former case, it is the opposite, and in the second case, there is little relation. Although this is not true, before the causes and treatments of TB and cancer were known, these views had enough popular support that they were useful as metaphors. This is unfortunate, as this would cause people to have views of people infected by tuberculosis or cancer which were not at all grounded in fact.

One point which I did not understand in the essay was that of the relationship of emotional sensitivity to tuberculosis. I do not understand where this relationship comes from, especially since observation, even in the eighteenth century, would seem to make this relationship seem awkward. Not all people with TB would be sensitive, caring people. But this view was treated as “scientific”. Today, we can at most call it metaphor, since we know that these are not facts. There is no causal relation between sensitivity and TB.

Besides using illness as a metaphor for certain types of individual behavior, the other use of illness as metaphor in society is to compare disagreeable societal conditions to disagreeable individual diseases, removing the individual from the disease. As disease (and death) frighten people, so would anything equated with “a cancer” or “a tumor”. Sontag claims that this comparison, like the personal comparisons above, is a “cheap shot”, that it is at best a limited comparison. Although I agree that use of terms like “cancer” to describe a problem causes an immediate response which may be unwarranted (such as Hitler’s describing of Jews as a cancer), this is not always the case. A cancer in a person is an unwanted growth, but one which is part of the body, diffcult to control or cure without causing damage to the carrier. A deeply rooted societal problem is very similar. A problem such as discrimination or racism is often so deeply rooted that it cannot be removed through normal law enforcement, since that would involve eliminating large segments of the population. It must be slowly treated, and it can bring about the downfall of a society.

After reading the essay, I see some of Sontag’s points, but I fail to see what direction the essay was going in. It seemed to lack a coherent thesis, but rather spent most of the time citing examples of metaphor in the literature, to what end I am not sure.